OpenStudy (anonymous):

help! Solve the system of equations algebraically. 4x-3y=8 5x-4y=5 a(16,21) b many solutions c no solution d(17,20)

OpenStudy (jordanloveangel):

Solution for 4x-3y=8 equation: Simplifying 4x + -3y = 8 Solving 4x + -3y = 8 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '3y' to each side of the equation. 4x + -3y + 3y = 8 + 3y Combine like terms: -3y + 3y = 0 4x + 0 = 8 + 3y 4x = 8 + 3y Divide each side by '4'. x = 2 + 0.75y Simplifying x = 2 + 0.75y

OpenStudy (jordanloveangel):

Solution for 5x-4y=5 equation: Simplifying 5x + -4y = 5 Solving 5x + -4y = 5 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '4y' to each side of the equation. 5x + -4y + 4y = 5 + 4y Combine like terms: -4y + 4y = 0 5x + 0 = 5 + 4y 5x = 5 + 4y Divide each side by '5'. x = 1 + 0.8y Simplifying x = 1 + 0.8y

OpenStudy (jordanloveangel):

so C